P/E and Growth Daisy Pixie Stix pays out 65% of its earnings as dividends. The firm has been earnings $0.10 cents per dollar of equity invested in the firm and investors require a 8.00% return. The last annual earnings were $2.00 per share. What is the P/E ratio of the stock?

28.56

36.65

14.44

23.11
Answer #1

C. 14.44

g = ROE x b = 0.10(1 – 0.65) = 3.50%

P/E = (1 – b) / (k – g) = 0.65 / (0.08 – 0.0350) = 14.44